### v1

# STeP: A General and Scalable Framework for Solving Video Inverse Problems with Spatiotemporal Diffusion Priors

Abstract
We study how to solve general Bayesian inverse problems involving videos using diffusion model priors. While it is desirable to use a video diffusion prior to effectively capture complex temporal relationships, due to the computational and data requirements of training such a model, prior work has instead relied on image diffusion priors on single frames combined with heuristics to enforce temporal consistency. However, these approaches struggle with faithfully recovering the underlying temporal relationships, particularly for tasks with high temporal uncertainty. In this paper, we demonstrate the feasibility of practical and accessible spatiotemporal diffusion priors by fine-tuning latent video diffusion models from pretrained image diffusion models using limited videos in specific domains. Leveraging this plug-and-play spatiotemporal diffusion prior, we introduce a general and scalable framework for solving video inverse problems. We then apply our framework to two challenging scientific video inverse problems—black hole imaging and dynamic MRI. Our framework enables the generation of diverse, high-fidelity video reconstructions that not only fit observations but also recover multi-modal solutions. By incorporating a spatiotemporal diffusion prior, we significantly improve our ability to capture complex temporal relationships in the data while also enhancing spatial fidelity. Our code is available at the GitHub repository STeP.
Figure 1: An overview of our proposed framework with Spatiotemporal Diffusion Priors (STeP) for video inverse problems. Left: We consider solving video inverse problems in scientific domains where a relatively large image dataset is available for training a prior but the amount of video data is limited. Middle: We propose a scalable and data-efficient spatiotemporal diffusion prior that directly models the video distribution using collections of both images and videos (samples generated from the prior are shown in the light purple box). We combine the prior and knowledge of the inverse problem in a state-of-the-art plug-and-play (PnP) diffusion solver [81]. Right: The resulting algorithm can recover multi-modal posterior distributions for difficult ill-posed inverse problems. Here we demonstrate the approach on the black hole video reconstruction problem, a highly nonlinear inverse problem with extremely sparse measurements leading to a multi-modal posterior. Specifically, we generated 100 posterior video samples and observe that there are three modes with equal data fidelity but with significantly different spatiotemporal structures (as shown by the frames and average optical flow from [59] in the last column). One of the recovered modes matches the ground truth in both spatial and temporal structure. This example shows the capability of our framework to generate diverse and high-fidelity video reconstructions for challenging scientific problems.
**footnotetext: These authors contributed equally to this work
1 Introduction
Using diffusion models as priors for solving Bayesian inverse problems has emerged as a powerful approach, demonstrating remarkable effectiveness in imposing image statistics learned from training data to guide recovered solutions. Plug-and-play (PnP) inversion methods that make use of diffusion priors have been successfully applied to diverse applications, including image restoration [67, 54, 34, 15, 48, 40, 87], medical imaging [30, 55, 13, 29, 18, 14, 81], and physics-based inverse problems [84, 1, 24, 56, 57, 73]. As a PnP prior, a diffusion model can be applied to various problems without retraining the model, making it flexible and easy to use. The success of these methods relies on two key factors: (1) access to a well-trained diffusion model, learned from a large set of unlabeled source data, and (2) a robust PnP framework capable of handling inverse problems with different underlying challenges [73, 81, 85, 84, 12].
Most prior work has focused on solving inverse problems for images; however, many critical inverse problems inherently involve temporal information, necessitating a general framework for solving video inverse problems [7, 65, 44, 79, 42]. Because training a video diffusion model is commonly believed to be too computationally challenging and data hungry, existing approaches to video inverse problems rely on image diffusion priors [17, 37, 36, 77], which process each frame independently together with various heuristics based on correlated noise or optical flow information to enforce temporal consistency (see Fig. 2 for a schematic illustration). However, these methods struggle to faithfully recover complex temporal relationships when observations become sparse and ill-posed, which is common in scientific inverse problems [65, 44].
In this paper, we propose a general and scalable approach for addressing video inverse problems, STeP, by integrating a SpatioTemporal video diffusion Prior into a PnP inversion method. To do so, we first demonstrate the feasibility of training a video diffusion prior for solving inverse problems using a limited amount of video data. Instead of training a video diffusion model from scratch, inspired by [66], we start from an image diffusion model and fine-tune the temporal modules, transforming it into a spatiotemporal video diffusion model using only a few hundred to a few thousand videos. This approach enables video diffusion in a data-efficient manner, drastically reducing training cost and making it feasible to obtain a video diffusion model from an image diffusion model within just a few hours on a single A100 GPU. After obtaining a well-trained spatiotemporal diffusion prior, we integrate it with a state-of-the-art PnP inversion method, namely the Decoupled Annealing Posterior Sampling (DAPS) [81] framework. STeP inherits the ability of DAPS to handle general inverse problems (with nonlinear forward models) and does not require additional temporal heuristics for solving video inverse problems.
We demonstrate the effectiveness of STeP on two challenging scientific video inverse problems: black hole video reconstruction (previewed in Fig. 1) and dynamic magnetic resonance imaging (MRI). Our experiments show that a fine-tuned spatiotemporal diffusion prior can be seamlessly integrated with the existing PnP diffusion solver, enabling efficient posterior sampling. As Fig. 1 illustrates, STeP not only achieves state-of-the-art results with improved temporal and spatial consistency but also effectively captures the multi-modal nature of highly ill-posed problems, recovering diverse plausible solutions from the posterior distribution. Notably, it achieves substantial improvements in temporal consistency, with a 6.50dB and 2.69dB increase in d-PSNR (average PSNR of difference images between all consecutive frames of a video) for black hole imaging and dynamic MRI, respectively—where d-PSNR quantifies temporal coherence. STeP also outperforms baselines in terms of spatial consistency by 1.69dB and 1.15dB in average frame-wise PSNR for black hole imaging and dynamic MRI, respectively.
Figure 2: A schematic comparison between prior works (top) and our STeP framework (bottom) for video inverse problems. The bold texts highlight the key differences between them. While prior works only use an image diffusion model and enforce temporal consistency with various heuristics, we directly learn a spatiotemporal diffusion prior. By leveraging a spatiotemporal prior, we improve both the temporal consistency and per-frame spatial consistency of the generated videos within a general and scalable PnP diffusion framework.
2 Background
2.1 Video latent diffusion models
Diffusion models [51, 53, 27, 54, 31] generate data by reversing a predefined noising process. Starting from the data distribution $p({\mathbf{x}}_{0})$, noisy data distributions $p({\mathbf{x}}_{t};\sigma_{t})$ are created by adding Gaussian noise with standard deviation $\sigma_{t}$, where $\sigma_{t}$ is a predefined noise schedule. To sample from the diffusion model, one requires the time-dependent score function $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})$ [51, 54, 31], which can be approximated by training a network ${\mathbf{s}}_{\bm{\theta}}({\mathbf{x}}_{t},\sigma_{t})$ using denoising score matching [64] with either a UNet [27, 32] or a transformer [43, 76] architecture.
While training diffusion models on the original high-dimensional data space may suffer from high computational cost, latent diffusion models (LDM) [46] instead generate an efficient, low-dimensional latent representation ${\mathbf{z}}_{0}$ of data ${\mathbf{x}}_{0}$ with a pretrained perceptual compression encoder $\mathcal{E}$ and decoder $\mathcal{D}$, which satisfy ${\mathbf{z}}_{0}=\mathcal{E}({\mathbf{x}}_{0})$ and $\mathcal{D}({\mathbf{z}}_{0})\approx{\mathbf{x}}_{0}$. The compression models $\mathcal{E}$ and $\mathcal{D}$ can be trained with VAE variants [35, 45, 8] with KL divergence regularization or VQGAN variants [26, 21, 63] with quantization regularization.
Video latent diffusion models are commonly believed to be hard to train due to computational cost in 3D modules in architecture and the requirement of a large video dataset [6, 86, 83, 28, 5]. Many recent methods tend to solve video modeling by training or fine-tuning from a pretrained image diffusion model with a video dataset.
2.2 Inverse problems with diffusion priors
Various methods have been proposed to solve Bayesian inversion using a pretrained diffusion model as a prior. Guidance-based methods [15, 47, 34, 50] approximate the intractable noisy likelihood score term $\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{y}}|{\mathbf{x}}_{t})$ while solving the reverse diffusion process. Variable splitting methods [16, 39, 49, 74, 73, 81] decompose inference into two alternating steps: one for enforcing the prior and another for incorporating the likelihood. Variational Bayes approaches [41, 24, 23] introduce a parameterized distribution to directly learn the posterior $p({\mathbf{x}}_{0}|{\mathbf{y}})$ with a diffusion prior. Sequential Monte Carlo (SMC) methods [71, 60, 10, 19] integrate diffusion sampling with SMC techniques to provide asymptotic convergence guarantees.
These methods rely on different requirements for the inverse problem and the prior. Some are specifically designed for linear problems [19, 37, 47], while others can handle nonlinear ones [73, 2]. Certain approaches require specialized designs for latent diffusion models [47, 49], whereas others can be naturally applicable [81, 39]. These requirements influence their generalizability and scalability for scientific inverse problems in different domains.
2.3 Video Inverse Problems (VIPs)
Recently, several works have extended diffusion-based approaches to video inverse problems (VIPs) [17, 88, 36, 37]. Some methods rely on image diffusion models with manually designed strategies, such as batch-consistent sampling (BCS) [37, 36] or using optical flow estimated from observations to warp the noise [17, 77]. However, we show that BCS has limited ability to faithfully recover the underlying temporal relationships for scientific inverse problems, in which the types of measurements considered often highly corrupt the temporal information. By implicitly imposing a static temporal prior via correlated noise, BCS struggles to recover complex temporal dynamics from such corrupted measurements. Previous work also highlighted a key limitation of optical flow-based methods [17, 77] being the reliance on accurate estimation of the optical flow from measurement data. When such measurements are sparse and contain limited temporal information, these methods struggle to impose an accurate temporal prior, leading to suboptimal reconstruction. Another line of work fine-tunes image diffusion models on domain-specific datasets to improve temporal consistency [88].
Compared to image diffusion models, data-driven video diffusion models offer a more general spatiotemporal prior for solving VIPs. In this work, we show the feasibility of training video diffusion models in scientific domains and their effectiveness in tackling video inverse problems.
3 Method
In this section, we introduce our framework, STeP, for solving video inverse problems with SpatioTemporal diffusion Priors. A schematic of our framework is provided in Fig. 2. We start by introducing our problem formulation in Sec. 3.1. We then propose a scalable and data-efficient way of training spatiotemporal diffusion priors in Sec. 3.2. We finally show in Sec. 3.3 that once such a prior is trained, it can be used to solve general video inverse problems.
Notations.
We adopt the following notations throughout the rest of the paper to avoid confusion. We use the variable ${\mathbf{x}}$ to denote objects in the image/video space and variable ${\mathbf{z}}$ to denote latent codes in the latent space. The variable ${\mathbf{y}}$ is always used for the measurements. Subscript $(\cdot)_{t}$ is the time index in the context of the diffusion process, where $t=0$ indicates the clean image. Superscript $(\cdot)^{[j]}$ is the index for the $j$-th frame in a video.
3.1 Basic formulation
We consider general video inverse problems (VIPs) of recovering an underlying target ${\mathbf{x}}_{0}$ from the measurements
${\mathbf{y}}=\mathcal{A}({\mathbf{x}}_{0})+{\mathbf{n}}$ (1)
where $\mathcal{A}(\cdot)$ is the forward model and ${\mathbf{n}}$ is the measurement noise. Importantly, ${\mathbf{x}}_{0}$ evolves over time, and substantial spatiotemporal information may be lost in the measurement process due to the ill-posed nature of $\mathcal{A}(\cdot)$, making it necessary to impose a prior on both the spatial and temporal dimensions of ${\mathbf{x}}_{0}$ for meaningful recovery. Our goal is to draw samples from the posterior distribution $p({\mathbf{x}}_{0}|{\mathbf{y}})\,\propto\,p({\mathbf{y}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})$. While the likelihood $p({\mathbf{y}}|{\mathbf{x}}_{0})$ can be derived from Eq. 1, it is often challenging to characterize the prior distribution $p({\mathbf{x}}_{0})$ for videos because of their high dimensionality and potentially limited number of samples for training.
Solve video inverse problems in latent space.
To overcome the challenge of high dimensionality, we propose to impose a spatiotemporal prior in latent space. Assuming that the set of likely ${\mathbf{x}}_{0}$’s is in the range of a decoder $\mathcal{D}$, we have that $\exists\,{\mathbf{z}}_{0}$ s.t. ${\mathbf{x}}_{0}=\mathcal{D}({\mathbf{z}}_{0})$ and can thus rewrite Eq. 1 as:
${\mathbf{y}}=\mathcal{A}(\mathcal{D}({\mathbf{z}}_{0}))+{\mathbf{n}}.$ (2)
It follows that the posterior $p({\mathbf{x}}_{0}|{\mathbf{y}})$ is the pushforward of the latent posterior $p({\mathbf{z}}_{0}|{\mathbf{y}})$ by $\mathcal{D}$, so it suffices to first generate latent samples from $p({\mathbf{z}}_{0}|{\mathbf{y}})$ and then decode them by $\mathcal{D}$.
3.2 Spatiotemporal diffusion prior
In order to meet the challenges of real-world VIPs, we aim for spatiotemporal diffusion priors with the following three properties:
  • (P1) It should be able to model distributions of high-resolution multi-frame videos and be reasonably efficient so that repeatedly calling it in a downstream solver would be computationally tractable.
  • (P2) It should directly learn temporal information from data instead of relying on heuristics so that it can capture sophisticated temporal dynamics and relationships.
  • (P3) It should be able to learn spatial information from both videos and images, given that static images are usually much more abundant for training than videos.
Our design of spatiotemporal diffusion priors aligns with each of these three properties, as discussed below.
Figure 3: Architecture of the spatiotemporal module in the proposed spatiotemporal UNet. Given a pretrained image diffusion UNet, we incorporate a zero-initialized temporal module with an ON/OFF switch into each 2D spatial module and initialize the additive weights ($\alpha$) to zero. Thus, it will add no effect at the start of fine-tuning and gradually learn by the video training data. The number of frames, height, and width are denoted by $n_{f}$, $n_{h}$, and $n_{w}$, respectively. The numbers of channels for input features ($f_{\text{in}}$) and output features ($f_{\text{out}}$) are denoted by $n_{\text{in}}$ and $n_{\text{out}}$, respectively.
Latent diffusion model with image encoder (P1).
We start by training a VAE [35] using the standard $L_{1}$ reconstruction loss with a scaled KL divergence loss on an image dataset. The KL divergence scaling factor is set to much less than 1 to prevent excessive regularization of the latent space. This allows us to obtain an image encoder $\mathcal{E}$ and decoder $\mathcal{D}$. Once they are trained, we fix their parameters and train a 2D UNet model ${\mathbf{s}}_{{\bm{\theta}}}({\mathbf{z}}_{t};\sigma_{t})$ using the standard denoising score matching loss. Despite recent progress in 3D spatiotemporal encoders and decoders [75, 72, 11], we opt for a 2D spatial encoder and decoder that processes each frame independently. This choice is due to efficiency considerations for the downstream PnP diffusion solver, where the decoder $\mathcal{D}$ is called multiple times during posterior sampling.
Spatiotemporal UNet as score function (P2).
Leveraging recent advancements in video generation [66, 28], we use a spatiotemporal UNet architecture to parameterize the time-dependent video score function, i.e. ${\mathbf{s}}_{{\bm{\theta}}}({\mathbf{z}}_{t};\sigma_{t})\approx\nabla_{{\mathbf{z}}_{t}}\log p({\mathbf{z}}_{t};\sigma_{t})$. The key component in the architecture is a spatiotemporal module for 3D modeling, as illustrated in Fig. 3. Given a pretrained image diffusion UNet, we introduce a zero-initialized temporal module for each 2D spatial module. Specifically, for an input feature $f_{\text{in}}$, let $f_{\text{out}}$ be the output of the spatiotemporal module, with $f_{\text{spat}}$ and $f_{\text{temp}}$ representing the outputs of the spatial and temporal branches, respectively. These features are combined using an alpha blending mechanism:
$f_{\text{out}}=(1-\alpha)\cdot f_{\text{spat}}+\alpha\cdot f_{\text{temp}},$ (3)
where $\alpha\in\mathbb{R}$ is a learnable parameter initialized as 00 in each spatiotemporal module. This design allows us to inherit the weights of the 2D spatial modules from the pretrained image diffusion model, significantly reducing the required training time. Additionally, by factorizing the 3D module into a 2D spatial module and a 1D temporal module, the spatiotemporal UNet only has marginal computational overhead compared to the original 2D UNet, striking a good balance between model capacity and efficiency.
Image-video joint fine-tuning (P3).
For compatibility with both image and video inputs, we introduce an ON/OFF switch signal in the spatiotemporal module. When the switch is set to OFF (indicating image input), the temporal module is disabled (or equivalently set $\alpha=0$). This ensures the output to $f_{\text{out}}=f_{\text{spat}}$ and reduces the spatiotemporal module to the original 2D spatial module, which processes each frame independently. During training, we initialize the weights of the spatial modules based on a pretrained image diffusion model and fine-tune all parameters of the spatiotemporal UNet using both image and video data. During fine-tuning, the model receives video data with probability $p_{\text{joint}}\in[0,1]$ and receives a pseudo video, where each frame is randomly sampled from an image dataset, with probability $1-p_{\text{joint}}$. The probability $p_{\text{joint}}$ is a tunable hyperparameter controlling the proportion of real video data in training. Pseudo video regularization helps the spatiotemporal UNet retain the spatial capabilities of the initialized spatial UNet. This strategy stabilizes training and prevents overfitting to the video dataset, proven effective in previous work [66].
3.3 Decoupled annealing posterior sampling
After obtaining a spatiotemporal diffusion prior, it is theoretically possible to combine it with any PnP diffusion solver. In this work, we employ the Decoupled Annealing Posterior Sampling (DAPS) framework, which is a novel framework for solving general inverse problems [81]. It is also easily compatible with latent diffusion models, making it an ideal choice for our purpose.
The core idea of DAPS is to sample the target latent posterior $p({\mathbf{z}}_{0}|{\mathbf{y}})$ by sequentially sampling $p({\mathbf{z}}_{t}|{\mathbf{y}})$ from $t=T$ to $t=0$. To do so, DAPS starts from $p({\mathbf{z}}_{T}|{\mathbf{y}})\approx\mathcal{N}({\bm{0}},\sigma_{\max}^{2}{\bm{I}})$ and sequentially draws a sample from $p({\mathbf{z}}_{t_{i-1}}|{\mathbf{y}})$ given a sample from $p({\mathbf{z}}_{t_{i}}|{\mathbf{y}})$ for $i=N,...,1$ based on a time schedule $\{t_{i}\}_{i=1}^{N}$. As shown by Proposition 1 of [81], this is possible if one can sample from:
$p({\mathbf{z}}_{0}|{\mathbf{z}}_{t},{\mathbf{y}})=\frac{p({\mathbf{y}}|{\mathbf{z}}_{0},{\mathbf{z}}_{t})p({\mathbf{z}}_{0}|{\mathbf{z}}_{t})}{p({\mathbf{y}}|{\mathbf{z}}_{t})}\,\propto\,p({\mathbf{y}}|{\mathbf{z}}_{0})p({\mathbf{z}}_{0}|{\mathbf{z}}_{t}).$
Indeed, by accessing the gradient, this unnormalized distribution can be sampled by MCMC methods, such as Langevin Monte Carlo (LMC) [69] and Hamiltonian Monte Carlo (HMC) [3]. After obtaining ${\hat{\mathbf{z}}_{0}}\sim p({\mathbf{z}}_{0}|{\mathbf{z}}_{t_{i}},{\mathbf{y}})$, we can easily sample from $p({\mathbf{z}}_{t_{i-1}}|{\mathbf{y}})$ by sampling ${\mathbf{z}}_{t_{i-1}}\sim\mathcal{N}({\hat{\mathbf{z}}_{0}},\sigma_{t_{i-1}}^{2}{\bm{I}})$ due to Proposition 1 of [81]. The pseudocode and more technical details of the proposed algorithm are provided in Appendix A.
4 Experiments
Table 1: Results on Black Hole Imaging for a Test Dataset of 20 Videos. We report the average with the standard deviation in parentheses. We use d-PSNR and d-SSIM to refer to PSNR and SSIM computed over consecutive frames. Due to the high ill-posedness of black hole imaging, we select the best out of five i.i.d. posterior samples based on the lowest average $\tilde{\bm{\chi}}_{\text{cp}}^{2}$ and $\tilde{\bm{\chi}}_{\text{logca}}^{2}$ for each test video. We find that the proposed spatiotemporal prior significantly enhances the temporal consistency (see middle columns) and improves per-frame spatial consistency (see left columns) compared to the baselines. As a result, the sampled video reconstructions better align with the observations, as shown in the better data fitting $\tilde{\bm{\chi}}^{2}$ statistics.
Methods PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) d-PSNR ($\uparrow$) d-SSIM ($\uparrow$) FVD ($\downarrow$) $\tilde{\bm{\chi}}_{\text{cp}}^{2}$ ($\downarrow$) $\tilde{\bm{\chi}}_{\text{logca}}^{2}$ ($\downarrow$)
STeP (ours) 27.23 (3.26) 0.75 (0.12) 0.172 (0.077) 39.05 (4.26) 0.95 (0.04) 192.34 1.907 (1.422) 1.403 (0.589)
IDM+BCS 25.54 (2.44) 0.74 (0.09) 0.183 (0.051) 32.54 (3.99) 0.94 (0.02) 255.41 2.411 (0.219) 2.380 (0.767)
IDM 24.13 (2.30) 0.69 (0.10) 0.196 (0.061) 29.42 (2.30) 0.92 (0.05) 1336.23 3.483 (3.454) 2.789 (2.753)
Figure 4: Visual results of black hole imaging. Left: We display selected frames from the ground truth and sampled video reconstructions from each method, with a stride of 8 frames. STeP achieves the best visual quality, accurately recovering the ring radius, bright spot location, and overall black hole appearance. Right: We visualize temporal dynamics by averaging the delta between consecutive frames over 4, 16, and 64 frames, respectively. STeP faithfully recovers motion patterns that align with the ground truth. In contrast, IDM+BCS underestimates the degree of motion and IDM lacks a consistent motion pattern. These spatiotemporal structures are further evident in the averaged optical flow over the entire 64 frames, where the optical flow is estimated using a pretrained model from [59].
We demonstrate the effectiveness of STeP on two challenging scientific inverse problems: black hole imaging [22] (Sec. 4.2) and dynamic MRI [25] (Sec. 4.3). We also provide an ablation study on the effectiveness of the image-video joint fine-tuning technique in Sec. 4.4. We provide additional experimental results and visualizations in Appendix E.
4.1 Baselines & Metrics
Baselines.
We establish a comparison by introducing two baselines. The first baseline replaces the video diffusion prior with an image diffusion prior, which is applied independently to each frame (referred to as IDM). The second baseline leverages the batch-consistency sampling technique [37] with an image diffusion prior (referred to as IDM+BCS). While IDM treats each frame as an independent image inverse problem, IDM+BCS enforces temporal consistency implicitly via correlated noise.
Metrics.
We evaluate frame-wise similarity between generated and ground truth videos using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [68], and Learned Perceptual Image Patch Similarity (LPIPS) [82]. These metrics are computed independently for each frame and then averaged. We use the versions implemented in piq [33] with all images normalized to the range $[0,1]$. For gray-scale frames, we repeat them 3$\times$ along the channel dimension before calculating the LPIPS score.
To assess temporal consistency, we introduce d-PSNR and d-SSIM, which compute PSNR and SSIM over the delta between consecutive frames. These metrics are also averaged across all delta frames. Additionally, we compute the Fréchet Video Distance (FVD) [61] between the test dataset and all video reconstructions to measure distributional similarity.111 We use the following project to compute FVD: https://github.com/JunyaoHu/common_metrics_on_video_quality.
Finally, we report the measurement data consistency using domain-specific metrics. For dynamic MRI, we report the mean squared error $\|\mathcal{A}({\mathbf{x}})-{\mathbf{y}}\|_{2}$ as data misfit. For black hole imaging, we use the $\bm{\chi}^{2}$ statistic (referred to Eq. 16 for detailed definition) on two closure quantities: the closure phase ($\bm{\chi}_{\text{cp}}^{2}$) and log closure amplitude ($\bm{\chi}_{\text{logca}}^{2}$). A $\bm{\chi}^{2}$ value close to 1 indicates good data fitting (refer to Appendix B for more detail). To facilitate a comparison between underfitting $(\bm{\chi}^{2}>1)$ and overfitting $(\bm{\chi}^{2}<1)$, we report a unified metric defined as:
$\tilde{\bm{\chi}}^{2}=\bm{\chi}^{2}\cdot\mathds{1}\{\bm{\chi}^{2}\geq 1\}+\frac{1}{\bm{\chi}^{2}}\cdot\mathds{1}\{\bm{\chi}^{2}<1\}.$ (4)
Table 2: Results on Dynamic MRI with 6$\times$ acceleration for a Test Dataset of 20 Videos. We report average PSNR, SSIM of the real and imaginary components and do similarly for d-PSNR, d-SSIM. The standard deviations are included in parentheses. LPIPS and FVD scores are calculated over the complex amplitude. The results show that by leveraging the proposed spatiotemporal prior, STeP consistently improves both temporal and per-frame spatial consistency.
Methods PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) d-PSNR ($\uparrow$) d-SSIM ($\uparrow$) FVD ($\downarrow$) Data Misfit
STeP (ours) 38.85 (1.50) 0.96 (0.01) 0.089 (0.019) 45.61 (2.45) 0.98 (0.01) 2153.34 10.31 (0.98)
IDM+BCS 37.51 (1.24) 0.95 (0.01) 0.095 (0.018) 42.92 (1.82) 0.96 (0.01) 2683.83 10.63 (0.94)
IDM 37.70 (0.99) 0.95 (0.01) 0.095 (0.018) 42.73 (1.65) 0.96 (0.01) 2789.80 10.61 (0.94)
Figure 5: Visual results of dynamic MRI. Left: We visualize the complex amplitude of a few selected frames of the ground truth and generated video reconstructions, including optical flows and zoom-ins. Each image is scaled according to the 95-th percentile calculated from the video dataset (i.e. around $95\%$ values are within the range $[0,1]$). STeP provides more accurate reconstructions with better main structure (the valve in the pink box) and fine-grained details (the light-color tissue in the green box). Right: Similarly to the black hole results, we show the temporal dynamics by visualizing the averaged delta between consecutive frames over 4 frames and 12 frames and the average optical flow estimated with TV-L1 optical flow [78]. STeP matches most of the motion patterns in the ground truth.
4.2 Black hole video reconstruction
Problem setup.
The goal is to reconstruct a video ${\mathbf{x}}_{0}\in\mathbb{R}^{n_{f}\times n_{h}\times n_{w}}$ of a rapidly moving black hole. Each measurement, or visibility, is given by correlating the measurements from a pair of telescopes to sample a particular spatial Fourier frequency of the source with very long baseline interferometry (VLBI) [62, 80]. Mathematically, the measured visibility given by the telescope pair $\{a,b\}$ for the $j$-th frame is:
${\mathbf{V}}_{\{a,b\}}^{[j]}=g_{a}^{[j]}g_{b}^{[j]}e^{-i\left(\phi_{a}^{[j]}-\phi_{b}^{[j]}\right)}{\mathbf{I}}_{\{a,b\}}^{[j]}({\mathbf{x}}_{0})+{\mathbf{n}}_{\{a,b\}}^{[j]},$ (5)
where ${\mathbf{I}}_{\{a,b\}}^{[j]}({\mathbf{x}}_{0})\in\mathbb{C}$ is the corresponding ideal visibility. Notably, ${\mathbf{V}}_{\{a,b\}}^{[j]}$ is a corrupted version of ${\mathbf{I}}_{\{a,b\}}^{[j]}({\mathbf{x}}_{0})$ that experiences Gaussian thermal noise ${\mathbf{n}}_{\{a,b\}}^{[j]}$ as well as telescope-dependent amplitude errors $g_{a}^{[j]}$, $g_{b}^{[j]}$ and phase errors $\phi_{a}^{[j]}$, $\phi_{b}^{[j]}$ [20]. To mitigate the impact of these amplitude and phase errors, closure quantities are derived and used to constrain inference [4]. Specifically, closure phases and log closure amplitudes are considered and can be written as:
${\mathbf{y}}_{\text{cp},\{a,b,c\}}^{[j]}$ $=\angle\left({\mathbf{V}}_{\{a,b\}}^{[j]}{\mathbf{V}}_{\{b,c\}}^{[j]}{\mathbf{V}}_{\{a,c\}}^{[j]}\right)\in\mathbb{R},$ (6)
${\mathbf{y}}_{\text{logca},\{a,b,c,d\}}^{[j]}$ $=\log\left(\frac{\left|{\mathbf{V}}_{\{a,b\}}^{[j]}\right|\left|{\mathbf{V}}_{\{c,d\}}^{[j]}\right|}{\left|{\mathbf{V}}_{\{a,c\}}^{[j]}\right|\left|{\mathbf{V}}_{\{b,d\}}^{[j]}\right|}\right)\in\mathbb{R}.$ (7)
Here, ($\angle$($\cdot$)∠ ( ⋅ ) and ||$\cdot$|| ⋅ | denote the complex angle and amplitude. The overall forward model is a combination of the two groups of closure quantities and an additional flux constraint (see Appendix B for more details). The likelihood function $p({\mathbf{y}}\mid{\mathbf{x}}_{0})$ is given by Eq. 16.
Dataset & spatiotemporal prior.
Measurements are simulated under observational conditions similar to those of the real data currently available for black hole video reconstruction. Namely, the Event Horizon Telescope (EHT) array observed the black hole Sagittarius A* over the course of a night in 2017, with approximately 100 minutes of that observation used for video reconstruction in [22]. A dataset of simulated black hole videos is compiled to match the expected dynamics of Sagittarius A* over this timescale. Specifically, we consider general relativistic magnetohydrodynamic (GRMHD) simulations [70] of the Sagittarius A* black hole under different black hole model assumptions and viewing conditions. The entire dataset contains 648 black hole videos, each with 1000 frames at 400$\times$400 spatial resolution. We then downsampled to 64 frames at 256$\times$256 spatial resolution, so $n_{f}=64$. We adopt a 64×64$\times$64 × (8×8$\times$8 × for both height and width) compression encoder and decoder. Details of the training hyperparameters are shown in Tab. 4.
Results.
We show the quantitative results in Tab. 1, and qualitative comparisons in Fig. 4. More results are provided in Appendix E. Quantitative evaluation shows that STeP significantly outperforms the baseline in temporal consistency. Visualizing the averaged delta frames reveals that our generated videos exhibit motion patterns closely matching the ground truth. In contrast, IDM+BCS produces more static motion due to its implicit static temporal prior assumption, while IDM lacks temporal consistency, leading to high incoherence across individual frames. This inconsistency arises because measurements are extremely sparse per frame. By leveraging a spatiotemporal prior, STeP jointly fits measurements across the entire video.
Multi-modal posterior analysis.
As discussed earlier, black hole imaging is a non-convex and highly ill-posed problem with extremely sparse measurements. Our experimental results in Fig. 1 indicate that its posterior distribution can be multi-modal. This implies that generated samples may align with distinct modes that, while differing significantly from the true videos, still fit the measurement data well. These findings demonstrate that STeP can generate diverse yet equally plausible videos, which is desirable for scientific discovery and uncertainty quantification.
4.3 Dynamic MRI
Problem setup.
We consider a dynamic MRI reconstruction problem in cardiac imaging, where the objective is to recover a video ${\mathbf{x}}_{0}\in\mathbb{C}^{n_{f}\times n_{h}\times n_{w}}$ of the heart from the subsampled Fourier space (a.k.a $k$-space) measurements ${\mathbf{y}}$. Mathematically, this can be formulated as
${\mathbf{y}}^{[j]}={\mathbf{m}}^{[j]}\odot\mathcal{F}\left({\mathbf{x}}_{0}^{[j]}\right)+{\mathbf{n}}^{[j]}\in\mathbb{C}^{n}\quad\text{for}j=1,...,n_{f},$
where ${\mathbf{m}}^{[j]}\in\{0,1\}^{n_{h}\times n_{w}}$ is the subsampling mask for the $j$-th frame, $\odot$ denotes element-wise multiplication, $\mathcal{F}$ is the Fourier transform, and ${\mathbf{n}}^{[j]}$ is the measurement noise. In our experiments, we used subsampling masks with an equi-spaced pattern (similar to those visualized in [65]) of both 6×6$\times$6 × acceleration with 24 auto-calibration signal (ACS) lines (Tab. 2) and 8×8$\times$8 × acceleration with 12 ACS lines (Fig. 6). For dynamic MRI, we use the Gaussian likelihood function: $\log p({\mathbf{y}}|{\mathbf{x}}_{0})\propto-\|\mathcal{A}({\mathbf{x}}_{0})-{\mathbf{y}}\|^{2}_{2}$.
Dataset & spatiotemporal prior.
We use the publicly available cardiac cine dataset from the CMRxRecon Challenge 2023 [65]. The entire dataset contains 3,324 cardiac MRI sequences with fully sampled and ECG-triggered $k$-space data from 300 patients, including various canonical views in cardiac imaging. The cardiac cycle was segmented into 12 temporal states, making each scan a 2D video of 12 frames, i.e. $n_{f}=12$. Given the fully sampled $k$-space data, we obtain the target videos by taking the inverse Fourier transform and resize all videos to the same spatial dimension of 192$\times$192. The measurements were generated by retrospectively applying the subsampling mask $\{{\mathbf{m}}^{[j]}\}_{j=1}^{n_{f}}$ to the fully sampled $k$-space data. We adopt a 16×16$\times$16 × (4×4$\times$4 × for both height and width) compression encoder and decoder. The detailed training hyperparameters are shown in Tab. 4.
Results.
We present the quantitative results for dynamic MRI in Tab. 2, with qualitative comparisons shown in Fig. 5. Additional results are provided in Appendix E. Unlike the ill-posed and non-convex nature of black hole imaging, dynamic MRI is a linear inverse problem focused on recovering fine-grained details. To highlight this, we zoom in on relevant structures in both video frames and averaged delta frames visualizations. Quantitative evaluation further demonstrates that STeP significantly outperforms the baseline in temporal consistency, which also enhances spatial alignment in the generated videos. More results are shown in Appendix E.
4.4 Effectiveness of Image-Video Joint Training
To better understand the impact of the spatiotemporal prior on solving inverse problems, we evaluate results using various checkpoints of the spatiotemporal UNet, each representing a prior fine-tuned for a different number of epochs. We assess performance using PSNR (blue curve), d-PSNR (red curve), and a data-fitting metric (green curve), as shown in Fig. 6 with a shared horizontal axis indicating the fine-tuning epochs. Since the spatiotemporal UNet is initialized from a pretrained image diffusion model, these curves reveal the gradual enhancement as increasingly stronger spatiotemporal priors are incorporated. The results indicate that temporal consistency and spatial consistency improve in a steady, synchronized manner as the prior undergoes further fine-tuning, evidenced by the close alignment of the blue and red curves. Furthermore, a better spatiotemporal prior enhances data fitting, as shown by the downward trend of the green curve.
Figure 6: Consistent improvement in image-video joint fine-tuning. We evaluate intermediate checkpoints for solving the inverse problems of (a) black hole imaging and (b) dynamic MRI with 8$\times$ acceleration. Both spatial quality (measured by PSNR) and temporal consistency (measured by d-PSNR) show steady improvement. For black hole imaging, the average chi-square metric is computed as the mean of $\tilde{\bm{\chi}}_{\text{cp}}^{2}$ and $\tilde{\bm{\chi}}_{\text{logca}}^{2}$.
5 Conclusion
We introduced STeP, a general framework for solving video inverse problems (VIPs) with a spatiotemporal diffusion prior. We demonstrated that it is possible to efficiently train a diffusion prior for videos, even with limited video data, enabling seamless integration into an existing PnP method for video inversion. By capturing complex temporal structure in the diffusion prior, our approach eliminates the need for temporal heuristics and enables the recovery of intricate temporal dynamics that resemble those in the training videos. We applied our method to two challenging scientific VIPs—black hole imaging and dynamic MRI—where it outperformed existing approaches in recovering both fine-grained spatial details and underlying temporal relationships. These results highlight that, with our proposed strategy, a diffusion video prior can be leveraged in a straightforward manner to tackle complex video inverse problems.
Acknowledgement
This research is funded by NSF award 2048237 and NSF award 2034306. B.Z is supported by the Kortschak Scholars Fellowship. Z.W. is supported by the Amazon AI4Science fellowship. B.F. is supported by the Pritzker Award.
We thank Ben Prather, Abhishek Joshi, Vedant Dhruv, C.K. Chan, and Charles Gammie for the synthetic black hole images GRMHD dataset used here, generated under NSF grant AST 20-34306.
Supplementary Material
Appendix A Detailed implementation of STeP
Here, we summarize the proposed framework for solving video inverse problems in Algorithm 1.
Algorithm 1 STeP: a general and scalable framework for solving video inverse problems with SpatioTemporal Prior
1:Discretization time steps $\{t_{i}\}_{i=1}^{N}$ where $t_{0}=0$ and $t_{N}=T$, noise schedule $\sigma_{t}$, likelihood $p({\mathbf{y}}\mid\cdot)$ with measurements ${\mathbf{y}}$, HMC step size $\eta$ and damping factor $\gamma$, number of HMC updates $M$, pretrained latent score function ${\mathbf{s}}_{\bm{\theta}}({\mathbf{z}};\sigma)\approx\nabla_{\mathbf{z}}\log p\left({\mathbf{z}};\sigma\right)$ with image decoder $\mathcal{D}$.
2:${\mathbf{z}}_{t_{N}}\sim\mathcal{N}({\bm{0}},\sigma_{t_{N}}^{2}{\bm{I}})$$\triangleright$ Initialization
3:for $i=N,...,1$ do
4:     ${\hat{\mathbf{z}}_{0}}\leftarrow\text{{Backward}}({\mathbf{z}}_{t_{i}};{\mathbf{s}}_{\bm{\theta}})$triangleright Solve PF-ODE (8) backward from $t=t_{i}$ to $t=0$
5:     ${\mathbf{p}}\sim\mathcal{N}({\bm{0}},{\bm{I}})$
6:     for $j=1,...,M$ do
7:         $\bm{\epsilon}_{j}\sim\mathcal{N}({\bm{0}},{\bm{I}})$
8:         $({\hat{\mathbf{z}}_{0}},{\mathbf{p}})$ = Hamiltonian-Dynamics(${\hat{\mathbf{z}}_{0}}$, ${\mathbf{p}}$, $\bm{\epsilon}_{j}$; $\eta$, $\gamma$) $\triangleright$ HMC updates for data consistency
9:     end for
10:     ${\mathbf{z}}_{t_{i-1}}\sim\mathcal{N}({\hat{\mathbf{z}}_{0}},\sigma_{t_{i-1}}^{2}{\bm{I}})$triangleright Proceed to the next noise level at time $t=t_{i-1}$
11:end for
12:return $\mathcal{D}({\mathbf{z}}_{t_{0}})$triangleright Return the decoded image
The algorithm’s main loop alternates between three key steps: (1) solving the PF-ODE backward from $t=t_{i}$ to $t=0$ (line 3), (2) performing multi-step MCMC updates (lines 4–8), and (3) advancing to the next noise level (line 9). We will discuss each step in detail.
Solving PF-ODE backward from $\bm{t=t_{i}}$ to $\bm{t=0}$
The probability flow ordinary differential equation (PF-ODE) [31] of the diffusion model, given by Eq. 8, governs the continuous increase or reduction of noise in the image when moving forward or backward in time. Here, $\dot{\sigma}_{t}$ denotes the time derivative of $\sigma_{t}$, and $\nabla_{{\mathbf{z}}_{t}}\log p({\mathbf{z}}_{t};\sigma_{t})$ represents the time-dependent score function [51, 54].
$\mathrm{d}{\mathbf{z}}_{t}=-\dot{\sigma}_{t}\sigma_{t}\nabla_{{\mathbf{z}}_{t}}\log p({\mathbf{z}}_{t};\sigma_{t})\,\mathrm{d}t,$ (8)
Our goal is to solve the probability flow ODE (PF-ODE), as defined in Eq. 8, backward from $t=t_{i}$ to $t=0$, given the intermediate state ${\mathbf{z}}_{t_{i}}$ and the pretrained latent score function ${\mathbf{s}}_{\bm{\theta}}({\mathbf{z}};\sigma)\approx\nabla_{\mathbf{z}}\log p({\mathbf{z}};\sigma)$. Any ODE solver, such as Euler’s method or the fourth-order Runge-Kutta method (RK4) [9], can be used to solve this problem. Following previous conventions [81], we adopt a few-step Euler method for solving it efficiently.
Multi-step MCMC updates
Any MCMC samplers can be used, such as Langevin Dynamic Monte Carlo (LMC) and Hamiltonian Monte Carlo (HMC). For example, the LMC update with step size $\eta$ is
${\mathbf{z}}_{0}^{+}={\mathbf{z}}_{0}+\eta\nabla_{{\mathbf{z}}_{0}}\log p({\mathbf{y}}\mid\mathcal{D}({\mathbf{z}}_{0}))+\eta\nabla_{{\mathbf{z}}_{0}}\log p({\mathbf{z}}_{0}\mid{\mathbf{z}}_{t})+\sqrt{2\eta}\bm{\epsilon}.$
Note that the first gradient term can be computed with (2). The second gradient term, on the other hand, can be calculated by
$\nabla_{{\mathbf{z}}_{0}}\log p({\mathbf{z}}_{0}\mid{\mathbf{z}}_{t})=\nabla_{{\mathbf{z}}_{0}}\log p({\mathbf{z}}_{t}\mid{\mathbf{z}}_{0})+\nabla_{{\mathbf{z}}_{0}}\log p({\mathbf{z}}_{0})\approx\nabla_{{\mathbf{z}}_{0}}\log p({\mathbf{z}}_{t}\mid{\mathbf{z}}_{0})+{\bm{s}}_{\bm{\theta}}({\mathbf{z}}_{0},t_{\min}).$
This approximation holds for $t_{\min}\approx 0$, assuming that ${\mathbf{z}}_{0}$ lies close to the clean latent manifold [52]. To improve both convergence speed and approximation accuracy, the MCMC samplers are initialized with the solutions obtained from the previous PF-ODE step, leveraging its outputs as a warm start.
Note that during MCMC updates, the decoder $\mathcal{D}$ needs to be evaluated multiple times in the backward pass. To accelerate this process, we adopt Hamiltonian Monte Carlo (HMC), which typically requires fewer steps for convergence, thereby speeding up the algorithm. For each multi-step MCMC update, we introduce an additional momentum variable ${\mathbf{p}}$, initialized as $\mathcal{N}({\bm{0}},{\bm{I}})$. The Hamiltonian-Dynamics(${\mathbf{z}}_{0},{\mathbf{p}},\bm{\epsilon};\eta,\gamma$) update with step size $\eta$ and damping factor $\gamma$ is given by:
${\mathbf{p}}^{+}$ $=(1-\gamma\eta)\cdot{\mathbf{p}}+\eta\nabla_{{\mathbf{z}}_{0}}\log p({\mathbf{z}}_{0}\mid{\mathbf{z}}_{t})+\sqrt{2\gamma\eta}\bm{\epsilon}$ (9)
${\mathbf{z}}_{0}^{+}$ $={\mathbf{z}}_{0}+\eta{\mathbf{p}}^{+}$ (10)
Proceeding to next noise level
According to Proposition 1 in [81], one can obtain a sample ${\mathbf{z}}_{t_{i-1}}\sim p({\mathbf{z}}_{t_{i-1}}\mid{\mathbf{y}})$ by simply adding Gaussian noise from a sample $\hat{\mathbf{z}}_{0}\sim p({\mathbf{z}}_{0}\mid{\mathbf{z}}_{t_{i}},{\mathbf{y}})$, given ${\mathbf{z}}_{t_{i}}\sim p({\mathbf{z}}_{t_{i}}\mid{\mathbf{y}})$ from last step. Thus we solve the target posterior sampling by gradually sampling from the time-marginal posterior of diffusion trajectory. The full parameters STeP is summarized in Tab. 3. The HMC-related parameters are searched on a leave out validation dataset consisting of 3 videos that are different from the testing videos.
Table 3: Hyper-parameters of STeP for black hole imaging and dynamic MRI. We provide and group the hyper-parameters of Algorithm 1.
Hyper-parameters Black hole imaging Dynamic MRI
PF-ODE Related
number of steps $N_{\text{ode}}$ 20 20
scheduler $\sigma_{t}$ $t$ $t$
HMC Related
number of steps $M$ 60 53
scaling factor $1-\gamma\eta$ 0.00 0.83
step size square $\eta^{2}$ 1.2e-5 1.2e-3
observation noise level $\sigma_{{\mathbf{y}}}$ 0.02 0.01
Decoupled Annealing Related
number of steps $N$ 25 20
final time $T$ 100 100
discretization time $\{t_{i}\},i=1,\cdots,N$ $\left(\tfrac{N-i}{N}\cdot T^{\frac{1}{7}}\right)^{7}$ $\left(\tfrac{N-i}{N}\cdot T^{\frac{1}{7}}\right)^{7}$
Appendix B Experimental Details
B.1 Black hole imaging
We introduce the black hole imaging (BHI) problem in more details. In Very Long Baseline Interferometry (VLBI), the cross-correlation of the recorded scalar electric fields at two telescopes, known as the ideal visibility, is related to the ideal source image ${\mathbf{x}}_{0}$ through a 2D Fourier transform, as given by the van Cittert-Zernike theorem [62, 80]. Specifically, the ideal visibility of the $j$-th frame of the target video is
${\mathbf{I}}_{\{a,b\}}^{[j]}({\mathbf{x}}_{0}):=\int_{\rho}\int_{\delta}\exp\left(-i2\pi\left(u_{\{a,b\}}^{[j]}\rho+v_{\{a,b\}}^{[j]}\delta\right)\right){\mathbf{x}}_{0}^{[j]}(\rho,\delta)\mathrm{d}\rho\mathrm{d}\delta\in\mathbb{C},$ (11)
where $(\rho,\delta)$ denotes the angular coordinates of the source image, and $\left(u_{\{a,b\}}^{[j]},v_{\{a,b\}}^{[j]}\right)$ is the dimensionless baseline vector between two telescopes $\{a,b\}$, orthogonal to the source direction.
Due to atmospheric turbulence and instrumental calibration errors, the observed visibility is corrupted by gain error, phase error, and additive Gaussian thermal noise [20, 58]:
${\mathbf{V}}_{\{a,b\}}^{[j]}:=g_{a}^{[j]}g_{b}^{[j]}\exp\left(-i\left(\phi_{a}^{[j]}-\phi_{b}^{[j]}\right)\right){\mathbf{I}}_{\{a,b\}}^{[j]}\left({\mathbf{x}}_{0}\right)+{\mathbf{n}}_{\{a,b\}}^{[j]}\in\mathbb{C}.$ (12)
where gain errors are denoted by $g_{a}^{[j]},g_{b}^{[j]}$, phase errors are denoted by $\phi_{a}^{[j]},\phi_{b}^{[j]}$, and thermal noise is denoted by ${\mathbf{n}}_{\{a,b\}}^{[j]}$. While the phase of the observed visibility cannot be directly used due to phase errors, the product of three visibilities among any combination of three telescopes, known as the bispectrum, can be computed to retain useful information. Specifically, the phase of the bispectrum, termed the closure phase, effectively cancels out the phase errors in the observed visibilities. Similarly, a strategy can be employed to cancel out amplitude gain errors and extract information from the visibility amplitude [4]. Formally, these quantities are defined as
${\mathbf{y}}_{\text{cp},\{a,b,c\}}^{[j]}$ $:=\angle({\mathbf{V}}_{\{a,b\}}^{[j]}{\mathbf{V}}_{\{b,c\}}^{[j]}{\mathbf{V}}_{\{a,c\}}^{[j]})\in\mathbb{R},$ (13)
${\mathbf{y}}_{\text{logca},\{a,b,c,d\}}^{[j]}$ $:=\log\left(\frac{\left|{\mathbf{V}}_{\{a,b\}}^{[j]}\right|\left|{\mathbf{V}}_{\{c,d\}}^{[j]}\right|}{\left|{\mathbf{V}}_{\{a,c\}}^{[j]}\right|\left|{\mathbf{V}}_{\{b,d\}}^{[j]}\right|}\right)\in\mathbb{R}.$
Here, ($\angle$($\cdot$)∠ ( ⋅ ) denotes the complex angle, and ||$\cdot$|| ⋅ | computes the complex amplitude. For a total of $P$ telescopes, the number of closure phase measurements ${\mathbf{y}}_{\text{cp},\{a,b,c\}}^{[j]}$ at is $\frac{(P-1)(P-2)}{2}$, and the number of log closure amplitude measurements ${\mathbf{y}}_{\text{logca},\{a,b,c,d\}}^{[j]}$ is $\frac{P(P-3)}{2}$, after accounting for redundancy. Since closure quantities are nonlinear transformations of the visibilities, the black hole imaging problem is non-convex.
To aggregate data over different measurement times and telescope combinations, the forward model of black hole imaging for the $j$-th frame can be expressed as
${\mathbf{y}}^{[j]}:=\left[\mathcal{A}_{\text{cp}}^{[j]}({\mathbf{x}}_{0}),\mathcal{A}_{\text{logca}}^{[j]}({\mathbf{x}}_{0}),\mathcal{A}_{\text{flux}}^{[j]}({\mathbf{x}}_{0})\right]:=\left[{\mathbf{y}}_{\text{cp}}^{[j]},{\mathbf{y}}_{\text{logca}}^{[j]},{\mathbf{y}}_{\text{flux}}^{[j]}\right],$ (14)
where ${\mathbf{y}}_{\text{cp}}^{[j]}=\left[{\mathbf{y}}_{cp,\{a,b,c\}}^{[j]}\right]$ is the set of all closure phase measurements and ${\mathbf{y}}_{\text{cp}}^{[j]}=\left[{\mathbf{y}}_{logca,\{a,b,c,d\}}^{[j]}\right]$ is the set of all log closure amplitude measurements for $j$-th frame. The total flux of the at $j$-th frame, representing the DC component of the Fourier transform, is given by
${\mathbf{y}}_{\text{flux}}^{[j]}:=\int_{\rho}\int_{\delta}{\mathbf{x}}_{0}^{[j]}(\rho,\delta)\mathrm{d}\rho\mathrm{d}\delta.$ (15)
The overall data consistency is an aggregation over all frames and typically expressed using the $\bm{\chi}^{2}$ statistics
where $\sigma_{\text{cp}}$, $\sigma_{\text{logca}}$, and $\sigma_{\text{flux}}$ are the estimated standard deviations of the measured closure phase, log closure amplitude, and flux, respectively, and $\beta$ is a hyperparameter that controls the strength of the flux regularization, which is empirically determined.
Our BHI experiments are based on the simulation of observing the Sagittarius A black hole with the EHT 2017 array of eight radio telescopes over an observation period of $\approx$100 minutes. We refer the readers to Fig. 5 of [38] for a visualization of the measurement patterns in Fourier space over time. To show the difficulty of this black hole video reconstruction problem, we visualize the dirty images obtained by applying inverse Fourier transform to the ideal visibilities, assuming no measurement errors, in Fig. 7. One can see that substantial spatiotemporal information is lost during the measurement process, so obtaining high-quality reconstructions relies on the effectiveness of incorporating prior information in the reconstruction process.
Figure 7: The dirty images from the ideal visibilities. We use the standard implementation in EHT library to get dirty images for each selected frame.
B.2 Dynamic MRI
MRI is an important imaging technique for clinical diagnosis and biomedical research. Despite its many advantages, MRI is known to be slow because of the physical limitations of the data acquisition in $k$-space. This leads to low patient throughput and sensitivity to patient’s motion [65]. To accelerate the scan speed, instead of fully sampling $k$-space, the compressed subsampling MRI (CS-MRI) technique subsamples $k$-space with masks $\{{\mathbf{m}}^{[j]}\}_{j=1}^{n_{f}}$. In our experiments, the 6×6$\times$6 × acceleration setting with 24 ACS lines leads to $\approx 73\%$ scan time reduction, while the 8×8$\times$8 × acceleration setting with 12 ACS lines leads to $\approx 82\%$ scan time reduction. Fig. 8 visualizes the subsampling masks used in our experiments, where $k_{x},k_{y}$ indicate the frequency encoding and phase encoding directions, respectively. The same mask is applied to the sampling of each individual frame of all videos.
Figure 8: Subsampling masks of 6×6$\times$6 × (left) and 8×8$\times$8 × (right) accelerations used in dynamic MRI experiments. The white areas in the center indicate the auto-calibration (ACS) signals. The horizontal and vertical directions are the frequency ($k_{x}$) and phase ($k_{x}$) encoding directions, respectively. The same mask is applied to the sampling of each individual frame of all videos.
B.3 Baseline Implementations
To make sure we are doing a fair comparison, we implement our two baselines according to Algorithm 1 by modifying several lines. We show the detailed modification below.
IDM+BCS
Follwing [37, 36], we replace the 3D spatiotemporal i.i.d. Gaussian noise in Algorithm 1 to batch consistent Gaussian noise, which is a 3D noise with identical 2D i.i.d. Gaussian frames, as shown in Eq. 17. To implement batch consistency sampling with image diffusion model, we only change the initial nose ${\mathbf{z}}_{t_{N}}$ (line 1 in Algorithm 1) and ${\mathbf{z}}_{t_{i-1}}$ (line 9 in Algorithm 1) from adding 3D spatiotemporal i.i.d. Gaussian noise to batch consistent noise. Moreover, the video diffusion is replaced with an image diffusion model that processes each frame independently.
$\bm{\epsilon}_{\text{BC}}^{[j]}=\bm{\epsilon},\quad\quad\bm{\epsilon}\in\mathbb{R}^{n_{h}\times n_{w}},\forall j=1,2,\cdots,n_{f}$ (17)
IDM
This is by replacing the video diffusion to an image diffusion model that processes each frame independently while keeping the remaining parts changed.
Appendix C Training Details for Video Diffusion Prior
Table 4: Hyper-parameters of the spatiotemporal video diffusion model. We provide and group the hyper-parameters according to each components in the model. The model is trained with 1 NVIDIA A100-SCM4-80GB GPU.
Hyper-parameters Black hole imaging Dynamic MRI
Dataset Related
frames $n_{f}$ 64 12
resolution $n_{h}\times n_{w}$ 256$\times$256 192$\times$192
$N_{\text{image}}$ 50000 39888
$N_{\text{video}}$ 648 3324
VAE Training Related
latent channels 1 2
block channels [64, 128, 256, 256] [256, 512, 512]
down sampling factor 8 4
batch size 16 16
epochs 25 10
$\beta_{\text{KL}}$ 0.06 0.03
IDM Training Related
block channels [128, 256, 512, 512] [128, 256, 512, 512]
batch size 16 16
epochs 200 50
Joint Fine-tuning Related
$p_{\text{joint}}$ 0.8 0.8
epochs 500 300
Other Info
VAE parameters 14.8M 57.5M
diffusion model parameters 131.7M 131.7M
VAE training time 4.5h 8.9h
++ image diffusion model training time 5.5h 3.8h
joint fine-tuning time 13.7h 22.8h
In this section, we show the detail of getting a video diffusion prior on black hole imaging and dynamic MRI, and we summarize the training hyper-parameters in Tab. 4. We define $D_{\text{image}}$ and $D_{\text{video}}$ as the image and video datasets, containing $N_{\text{image}}$ and $N_{\text{video}}$ data points, respectively. The image dataset $D_{\text{image}}$ includes all individual frames from the video dataset $D_{\text{video}}$, along with additional large-scale image data to enhance generalization. For data augmentation, we apply random horizontal/vertical flipping and random zoom-in-and-out to improve robustness and diversity in training.
We first train the compression functions, the encoder $\mathcal{E}$ and decoder $\mathcal{D}$, on an image dataset. The training objective consists of an L1 reconstruction loss combined with a KL divergence term scaled by a factor $\beta_{KL}$. The loss function for training is as defined in Eq. 18. The Adam optimizer is used as the default optimizer throughout the paper. The loss function for training the variational autoencoder (VAE) is given by:
$\mathcal{L}_{\text{VAE}}=\mathbb{E}_{q_{\phi}({\mathbf{z}}_{0}|{\mathbf{x}}_{0}),{\mathbf{x}}_{0}\sim D_{\text{image}}}\left[\|\mathcal{D}({\mathbf{z}}_{0})-{\mathbf{x}}_{0}\|_{1}\right]+\beta_{\text{KL}}D_{\text{KL}}\big{(}q_{\phi}({\mathbf{z}}_{0}|{\mathbf{x}}_{0})\|p({\mathbf{z}}_{0})\big{)}$ (18)
where $p({\mathbf{z}}_{0})$ is the standard Gaussian $\mathcal{N}({\bm{0}},{\bm{I}})$ and $q_{\phi}({\mathbf{z}}_{0}|{\mathbf{x}}_{0})$ is the isotropic Gaussian distribution over ${\mathbf{z}}_{0}$ where the mean and standard deviation is given by $\mathcal{E}({\mathbf{x}}_{0})$. Next, we train the image diffusion UNet ${\mathbf{s}}_{{\bm{\theta}}}$ using the standard score-matching loss, as defined in Eq. 19, following [27, 54].
$\mathcal{L}_{\text{IDM}}=\mathbb{E}_{{\mathbf{z}}_{0}\sim q_{\phi}({\mathbf{z}}_{0}\mid{\mathbf{x}}_{0}),x_{0}\sim D_{\text{image}},\epsilon\sim\mathcal{N}({\bm{0}},{\bm{I}}),t\sim\mathcal{U}(0,1)}\left[\sigma^{2}_{t}\left\|{\mathbf{s}}_{{\bm{\theta}}}({\mathbf{x}}_{t},t)-\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})\right\|^{2}\right]$ (19)
After pertaining, the image diffusion UNet ${\mathbf{s}}_{{\bm{\theta}}}$ is then converted to a spatiotemporal UNet by adding zero-initialized temporal modules to 2D spatial modules and fine-tune jointly with video and image datasets. We use the same Eq. 19 without change ${\mathbf{x}}_{0}$ to video or pseudo video input and use the encoder to process each frame independently.
After pretraining, the image diffusion U-Net ${\mathbf{s}}_{{\bm{\theta}}}$ is transformed into a spatiotemporal UNet by integrating zero-initialized temporal modules into the existing 2D spatial modules. The model is then fine-tuned jointly using both video and image datasets. We use the same loss as in Eq. 19, by changing ${\mathbf{x}}_{0}$ to a video or a pseudo-video input. Each frame is independently processed using the encoder $\mathcal{E}$, ensuring that spatial representations remain aligned while temporal consistency is learned through the added temporal modules.
Appendix D Discussion
D.1 Sampling Efficiency
We discuss the sample efficiency in this section. The sampling time of STeP  depends on the total number of video diffusion model calling $N_{\text{vdm}}$ and total number of decoder, and its gradient calling $N_{\text{dec}}$. We summarize these parameters and sampling requirement in Tab. 5.
Table 5: The sampling requirement and number of function callings in STeP for two problems. The run time and memory is tested using 1 NVIDIA A 100-SCM4-80GB GPU.
$N_{\text{dec}}$ $N_{\text{vdm}}$ time (s) memory (GB)
Black hole imaging 1500 500 645 52
Dynamic MRI 1060 400 332 23
D.2 Limitations and Future Extension
Though STeP is a general and scalable framework for solving video inverse problem with spatiotemporal diffusion prior, the sampling cost of STeP is high due to the requirement of backpropagation through decoder $\mathcal{D}$ in MCMC updates in Algorithm 1. This is forcing us to balance between the capability of the decoder and its computational cost. We leave the exploration of performing MCMC updates in pixel space or other approaches to bypass calling decoder as future work.
Appendix E More Results & Visualization
Dynamic MRI with higher acceleration.
To access the capability of using spatiotemporal prior for solving more challenging inverse problems, we increase the acceleration times in Dynamic MRI, which makes the observation more sparse. The results are summarized in Tab. 6.
Table 6: Results on Dynamic MRI with 8$\times$ acceleration for a Test Dataset of 20 Videos. Compared to the 6× acceleration results in Tab. 2, STeP achieves a significantly larger performance improvement over the baselines, further highlighting the effectiveness of the spatiotemporal prior.
Methods PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) d-PSNR ($\uparrow$) d-SSIM ($\uparrow$) FVD ($\downarrow$) Data Misfit
STeP (ours) 35.31 (2.76) 0.91 (0.04) 0.100 (0.024) 43.36 (3.29) 0.96 (0.02) 2316.83 8.41 (0.80)
IDM+BCS 31.95 (1.79) 0.85 (0.04) 0.123 (0.021) 37.41 (2.08) 0.89 (0.03) 3549.81 8.87 (0.74)
IDM 32.09 (1.34) 0.85 (0.03) 0.121 (0.020) 36.58 (1.74) 0.88 (0.03) 3530.57 8.86 (0.76)
More visualizations
Here, we show the VAE reconstruction results in Fig. 9, unconditional samples in Fig. 10 and additional posterior samples in Fig. 11.
Figure 9: Visualization of VAE Reconstructions.
Figure 10: Visualization of video diffusion model unconditional samples. The videos are sampled by solving PF-ODE with 100 Euler’s steps.
Figure 11: Visualization of STeP posterior samples. The videos are sampled using the Algorithm 1.